OpenStudy (liam):

how can i solve 6+4 (r-2)=r+7

OpenStudy (nowhereman):

Pull all summands with r to the one side and all without to the other side.

OpenStudy (nowhereman):

It's the same with the other questions. Try not to ask many equal questions.

OpenStudy (liam):

okay but how do i do that?

OpenStudy (nowhereman):

You can add the same things on both sides.

OpenStudy (liam):

can someone walk me through it? if you dont mind?

OpenStudy (anonymous):

get rid of the parenthesis 6 + 4r - 8 = r + 7 combine like terms 4r - 2 = r + 7 subtract r on both sides 3r - 2 = 7 add 2 on both sides 3r = 9 divide r = 3

OpenStudy (liam):

thank so much

OpenStudy (anonymous):

cheers